Dec 17, 2015 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Update BCAL clusterizer to improve agreement between simulated and generated MC (Mark)
  2. Modify covariance matrix for FCAL and BCAL to correspond to the current resolution of the detectors (Matt).
  3. Change F250 simulation code to match firmware and add flexibility for new algorithms (Mike)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run Preparations
    1. FCAL Update
    2. BCAL Update
  4. Reconstruction and Simulation
    1. Energy Spectra from Muon Decays in BCAL (Part 2) (Andrei and Irina)
  5. Calibration
    1. Re-Calibration of BCAL with Cosmics (Andrei and Irina)
    2. Calibration Challenge
  6. Any other business


Attendees:Elton, David, Adesh, Tegan (JLab); Zisis, Noemi, Andrei (UofR); Matt (IU); Will, Mike (CMU); George (Athens)

  1. Announcements
    • CEBAF delivered 12.05 GeV to the tagger dump on Monday evening. Milestone!
    • CEBAF delivered 11 GeV at 70 microA to Hall A. Milestone!
    • RF separation to run Hall A and D in parallel succeeded with pulsed beam, but PSS system is presently preventing cw beam
    • Hall D will be working on trigger and DAQ. The evening is scheduled for accelerator to work on the fast feedback system for Hall D.
    • The goal (hope) is to be able to run RF-separated beam to A and D over the weekend. Time is getting short: beam goes off at 6 am Monday morning.
  2. Action Items
    • No significant updates
  3. Run Preparations
    1. FCAL Update
      • Turned FCAL voltages on. 3 channels are disabled because they do not hold voltage continuously
      • Matt and Adesh replaced 95 bases over the last month, mostly bases that were dropping voltages to zero intermittently.
      • The problem bases have been shipped back to IU for diagnosis and repair
      • There are some additional bases that are drawing excessive current of about 14.5 mA (about 3x normal). Their reported voltages are slightly lower than their set point, but are otherwise ok. Will need to decide whether they should be replaced before the spring run.
      • Investigation of problem bases by Paul Smith at IU show two main classes of problems: 1) resister divider in middle circuit have grime underneath and must be cleaned carefully. 2) high current develops due to failed component that can be found using IR light. The primary culprit creates a short in the multiplier chain.
      • Future maintenance: Matt suggests that we train JLab techs to fix the main identified problems on the bases. IU can continue to support the effort and trouble shoot difficult failures. Fixes may require a fume hood.
    2. BCAL Update
      • Log Entry 3368145
      • BCAL seems to be happy at 10 deg with dew points below -20 degC. Should consider optimum running for spring.
      • New additions are the scaler displays for all systems (including BCAL and FCAL). At the moment FADC rates when the DAQ is not running are dubious.
  4. Reconstruction and Simulation
    1. Energy Spectra from Muon Decays in BCAL (Part 2) (Andrei and Irina)
      • Andrei sent additional txt files with information to Elton and Will. Elton and Will had some questions that will be passed on offline. Elton requested the data from the MC plots for ease of superimposing on measured data. Also histograms of total energy in calorimeter would be useful.
  5. Calibration
    1. Re-Calibration of BCAL with Cosmics (Andrei and Irina)
      • Andrei has completed a new calibration of thru muons in the calorimeter by using the calorimeter to determine track lengths through the calibrated cells.
      • Data used were from Dec 2014 (new analysis) and compared to previous calibration using data from Fall 2014 (old analysis). Relatively good consistency obtained for valid channels. Bias setting was 1.2 V above breakdown.
      • New calibration covers essentially all channels (including layer 4) and gives a very uniform response across all cells.
      • Comparison with Will's constants from pi0 calibration shows some discrepancies, especially in layer 4.
      • Andrei gave the latest calibration constants to Will, who used them to plot the pi0 mass for the 2014 data.
      • With the new constants, the pi0 peak falls at about 129 MeV and the width is about 12 MeV.
      • Elton: Would be interesting to obtain thru muon calibration for data taken last week to compare with current settings.
      • Andrei: Suggested that discrepancies between thru muons and pi0 calibrations could be due to z-dependence. Will confirmed that the calibration was plotted vs z and systematic dependencies were observed. Will wait till spring to take more data and be able to study this in more detail.
    2. Analysis
      • Tegan has been working with Justin on eta' analysis and making progress, but nothing to report.
    3. Calibration Challenge
  6. Any other business
    • Next meeting of the year will be the BCAL reconstruction on Jan 7. This keeps the same sequence of meetings from this year into the next.